How many liters of pure alcohol must be added to 20L of a 40% solution to obtain a 60% solution?
Answer by mananth(16946)

percent ---------------- quantity
40 ---------------- 20
100 ---------------- x
60 ---------------- 20+x
......
40*20+100*x=60( 20+x)
800+100 x=1200+60x
100x-60x=1200-800
40x=400
/ 40
x=10 liters of 100 % alchohol